# Using Dynamic Safety Margins as Control Barrier Functions

### v2

@@ 3 From Dynamic Safety Margins to Control Barrier Functions
Based on the intuition that both CBFs and DSMs can be used to design add-on modules for constraint handling, the following theorem states that a dynamic safety margin is, in fact, a CBF for an augmented system consisting of the concatenation of $\boldsymbol{\mathbf{x}}\in\mathbb{X}\subset\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{v}}\in\mathbb{V}\subset\mathbb{R}^{l}$
@@
where $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{U}\times\mathbb{R}^{l}$ are the augmented inputs.
@@
If $\Delta:\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}\to\mathbb{R}$ is a DSM for $\pi$, then it is a CBF with respect to the augmented system (22).
Proof 3.8.
Given $\alpha\in\mathscr{K}_{\infty}$, consider $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$. By (19b), $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, which implies that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. Consider the augmented input $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{U}\times\mathbb{R}^{l}$ and note that, by (19c),
$\dot{\Delta}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}$ $=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}\big{(}f(\boldsymbol{\mathbf{x}})+g(\boldsymbol{\mathbf{x}})\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}+\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}0$
$=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$
$\geq 0$
$\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
It then follows that, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\times\mathbb{V}$ such that $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0$,
$\sup_{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{U}\times\mathbb{R}^{l}}\left[\dot{\Delta}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\right]\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$ (23)
Therefore, $\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is a CBF.
For DSMs, the sets $\mathcal{C}\subset\mathbb{X}\times\mathbb{V}$ and $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\subset\mathbb{U}\times\mathbb{R}^{l}$ are
$\mathcal{C}$ =$=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{D}_{\boldsymbol{\mathbf{v}}}\!\times\!\mathbb{V}\mid\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0\}\subset\mathcal{X}\times\mathcal{V},$ (24)
$\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ =$=\{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{U}\times\mathbb{R}^{l}\mid\dot{\Delta}+\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}\geq 0\}.$ (25)
A significant concern in the CBF literature arises when the term $L_{g}h(\boldsymbol{\mathbf{x}})=0$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. This means $h$ lacks uniform relative degree 1 over $\mathcal{C}$, and implies a loss of controllability that can be catastrophic when $L_{f}h(\boldsymbol{\mathbf{x}})<-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}$. Although many relevant works address this issue (e.g. [7, 8]), the following lemma shows that $L_{g}h(\boldsymbol{\mathbf{x}})=0$ is never a concern for DSM-based CBFs.
Lemma 3.9.
Let $\Delta$ be a DSM for $\pi$ and let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{C}$ be such that $\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}=0$. Then, $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
Proof 3.10.
By construction, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Thus,
$\dot{\Delta}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}$ $=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})+\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}\begin{bmatrix}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ 0\end{bmatrix}$
@@
$=\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
@@
Therefore, any admissible augmented input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{U}\times\mathbb{R}^{l}$ trivially satisfies the CBF condition.
@@
Recalling Theorem 3.7, it is possible to reformulate the CBF policy $\beta$ for the augmented system $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. To do so, let $\boldsymbol{\mathbf{r}}\in\mathbb{V}$ be a desired reference, let $\kappa:\mathbb{X}\times\mathbb{V}\to\mathbb{U}$ be a nominal control policy with desirable closed-loop performance, and let $\rho:\mathbb{V}\times\mathbb{V}\to\mathbb{R}^{l}$ be a navigation field. The following theorem formulates the CBF-based policy $\beta$ for DSMs and the augmented system (22).
@@
Let $\Delta$ be a DSM and let $\mathcal{C}$, given in (24), be compact. Assume also that 00 is a regular value of $\Delta$ and let $\eta>0$ be a constant scalar. If the optimization-based policy $\beta:\mathcal{C}\to\mathcal{U}\times\mathbb{R}^{l}$
$\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\!\underset{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})}{\operatorname{argmin}}\|\boldsymbol{\mathbf{u}}\!-\!\kappa(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\|^{2}\!+\eta\|\boldsymbol{\mathbf{w}}\!-\!\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\|^{2},$ (26)
is locally Lipschitz continuous, then the trajectory of the augmented system (22) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\mathcal{C}$.
Proof 3.12.
By Theorem 3.7, $\Delta$ is a CBF and, since the set $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\neq\emptyset$, compactness on $\mathcal{U}$ is not needed to apply Lemma 1. So, the map $\beta$ is well-defined. Local Lipschitzness of $\beta$ guarantees existence and uniqueness of the solution to the closed-loop system. Safety follows immediately because regularity of 00 in $\Delta$ and the condition $\dot{\Delta}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\geq-\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}$, ensure invariance of $\mathcal{C}\subset\mathcal{X}\times\mathcal{V}$ for the dynamics (22) with input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$.
@@
Although the set $\mathcal{K}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ is generally not convex, if the input constraint set $\mathcal{U}$ is polyhedral, the optimization problem (26) becomes a quadratic program (QP) for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{C}$, which can be solved efficiently.
@@
We provide some conditions that guarantee the CBF policy $\beta$ for DSM-based CBFs is locally Lipschitz continuous. Let us assume the input constraint set is polyhedral $\mathcal{U}=\{\boldsymbol{\mathbf{u}}\in\mathbb{U}\mid M\boldsymbol{\mathbf{u}}\leq\boldsymbol{\mathbf{b}}\}$. We begin by defining the following linear program (LP), which finds the “width” of the feasible set [33]
$\omega^{*}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=$ $\max_{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}},\omega)\in\mathbb{U}\times\mathbb{R}^{l}\times\mathbb{R}}\quad\omega$ (27a)
s.t. $\begin{bmatrix}A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\end{bmatrix}\begin{bmatrix}\boldsymbol{\mathbf{u}}\\ \boldsymbol{\mathbf{w}}\end{bmatrix}+\boldsymbol{\mathbf{1}}\omega\leq b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (27b)
where $\boldsymbol{\mathbf{1}}$ is a column vector of $1$’s,
$A(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}-\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&-\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\\ M&0\end{bmatrix},$
and
$b(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\begin{bmatrix}\alpha\big{(}\Delta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}+\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\\ \boldsymbol{\mathbf{b}}\end{bmatrix}.$

### v3

@@ 3 From Dynamic Safety Margins to Control Barrier Functions
This section explores the connections between DSMs and CBFs, formulates a safe, optimization-based control policy and provides conditions for this policy to be locally Lipschitz continuous.
3.1 DSMs are CBFs
Based on the intuition that both CBFs and DSMs can be used to design add-on modules for constraint handling, the following theorem states that, a dynamic safety margin is, in fact, a CBF for an augmented system consisting of the concatenation of $\boldsymbol{\mathbf{x}}\in\mathbb{R}^{n}$ and $\boldsymbol{\mathbf{v}}\in\mathbb{R}^{l}$
@@ 3.1 DSMs are CBFs
where $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\mathbb{R}^{m}\times\mathbb{R}^{l}$ are the augmented inputs.
@@
If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM for $\pi$, then $\Delta$ is a (vector-valued) CBF with respect to the augmented system (24).
Proof 3.7.
Define the set-valued maps
$\tilde{\mathcal{C}}_{i}(c)=\{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}:0\leq\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq c\},$ (25)
and the functions $\hat{\alpha}_{i}:[0,\infty)\to\mathbb{R}$ by
@@
$\hat{\alpha}_{i}(c)=-\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(c)}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (26)
where $i\in\{1,\ldots,n_{c}\}$. Note that for any $c\in[0,\infty)$, the set $\tilde{\mathcal{C}}_{i}(c)\subset\tilde{\mathcal{C}}$ is compact. It can also be shown that the set-valued maps $\mathcal{C}_{i}$ are both upper- and lower-hemicontinuous. Thus, it follows by Berge’s maximum theorem [29] that each $\hat{\alpha}_{i}$ is continuous on $[0,\infty)$. Moreover, it is nondecreasing because for any $c_{1}\leq c_{2}$, $\tilde{\mathcal{C}}_{i}(c_{1})\subset\tilde{\mathcal{C}}_{i}(c_{2})$. Next, note that for any $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)$, $\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=0$, and it follows by (22c) that
$\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\geq 0.$
Therefore,
$\hat{\alpha}_{i}(0)=-\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(0)}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\leq 0.$ (27)
There always exists $\alpha_{i}\in\mathscr{K}_{\infty}$ such that $\forall c\in[0,\infty)$, $\hat{\alpha}_{i}(c)\leq\alpha_{i}(c)$. Define $\alpha(c)=\max_{i\in\{1,\ldots,n_{c}\}}\alpha_{i}(c)$ and note that $\alpha\in\mathscr{K}_{\infty}$. Let $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ and $i\in\{1,\ldots,n_{c}\}$ be given and note that since $\boldsymbol{\mathbf{x}}\in\mathcal{X}_{\boldsymbol{\mathbf{v}}}$, it follows that $\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\mathcal{U}$. With this,
$\dot{\Delta}_{i}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}$ $=\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})+\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}0$
$\geq\inf_{(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}_{i}(\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}))}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f_{\pi}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$
$\geq-\hat{\alpha}_{i}\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}$
$\geq-\alpha_{i}\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}$
$\geq-\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
We have shown that $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\mathcal{U}\times\mathbb{R}^{l}$ is such that
$\min_{i\in\{1,\ldots,n_{c}\}}\left[\dot{\Delta}_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})+\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}\right]\geq 0.$ (28)
We conclude that the channels $\Delta_{i}$ have the control-sharing property, so $\Delta$ is a (vector-valued) CBF.
Let us define the set of viable inputs at $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for DSM-CBFs as
[…]
The following corollary states that DSM-CBFs satisfy (9), so $\mathcal{U}$ needs not be compact to conclude control invariance of $\tilde{\mathcal{C}}$.
Corollary 3.8.
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$. Then, there exists $\alpha\in\mathscr{K}_{\infty}$ such that $\Delta$ satisfies (9). Specifically, $\forall(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. Moreover, if 00 is a regular value of each $\Delta_{i}$, then the set $\tilde{\mathcal{C}}\subset\tilde{\mathcal{X}}\subset\mathcal{X}\times\mathcal{V}$ is control invariant.
Proof 3.9.
The first statement was already shown in the proof of Theorem 3.6. The control invariance of $\tilde{\mathcal{C}}$ is an immediate consequence of Theorem 2 when $n_{c}=1$. For the general case $n_{c}>1$, we refer the reader to [28, Proposition 4.3.7].
A significant concern in the CBF literature arises when the term $L_{g}h(\boldsymbol{\mathbf{x}})=0$ for some $\boldsymbol{\mathbf{x}}\in\mathcal{C}$. This means $h$ lacks uniform relative degree 1 over $\mathcal{C}$, and implies a loss of controllability that can be catastrophic when $L_{f}h(\boldsymbol{\mathbf{x}})<-\alpha\big{(}h(\boldsymbol{\mathbf{x}})\big{)}$. Although many relevant works address this issue (e.g. [7, 8]), and Corollary 3.8 already shows it is not a concern for DSM-CBFs, the following corollary shows an additional desirable property of DSM-CBFs under total controllability loss.
Corollary 3.10.
Let $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ be a DSM for $\pi$ and let $\alpha\in\mathscr{K}_{\infty}$ be such that (9) holds. If $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$ is such that
$\begin{bmatrix}\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})&\frac{\partial\Delta}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}=0,$ (30)
then, $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
Proof 3.11.
Note that since $\big{(}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, it follows that $\forall i\in\{1,\ldots,n_{c}\}$,
$\dot{\Delta}_{i}\big{(}\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}},\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),0\big{)}\!$ $=\!\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\!+\!\begin{bmatrix}\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}g(\boldsymbol{\mathbf{x}})\!&\!\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{v}}}\end{bmatrix}\!\begin{bmatrix}\pi(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\\ 0\end{bmatrix}$
$=\frac{\partial\Delta_{i}}{\partial\boldsymbol{\mathbf{x}}}f(\boldsymbol{\mathbf{x}})\geq-\alpha\big{(}\Delta_{i}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\big{)}.$
@@
Therefore, constraint (28) is trivially satisfied at $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ for any augmented input $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})$. We conclude that $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\mathcal{U}\times\mathbb{R}^{l}$.
3.2 DSM-CBF Optimization-based Policy
@@
Recalling Theorem 3.6, it is possible to reformulate the CBF policy $\beta$ for the augmented system $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$. To do so, let $\boldsymbol{\mathbf{r}}\in\mathbb{R}^{l}$ be a desired reference, let $\kappa:\mathbb{R}^{n}\times\mathbb{R}^{l}\to\mathbb{R}^{m}$ be a nominal control policy with desirable closed-loop performance, and let $\rho:\mathbb{R}^{l}\times\mathbb{R}^{l}\to\mathbb{R}^{l}$ be a navigation field. The following theorem formulates the CBF-based policy $\beta$ for DSMs and the augmented system (24).
@@
Let $\eta>0$ be a constant scalar. If $\Delta:\tilde{\mathcal{D}}\to\mathbb{R}^{n_{c}}$ is a DSM, then there exists $\alpha\in\mathscr{K}_{\infty}$ such that the optimization-based policy $\beta:\tilde{\mathcal{C}}\to\mathcal{U}\times\mathbb{R}^{l}$
$\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})=\!\underset{(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})\in\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})}{\operatorname{argmin}}\|\boldsymbol{\mathbf{u}}\!-\!\kappa(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{r}})\|^{2}\!+\eta\|\boldsymbol{\mathbf{w}}\!-\!\rho(\boldsymbol{\mathbf{v}},\boldsymbol{\mathbf{r}})\|^{2},$ (31)
is feasible for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$. Moreover, if $\beta$ is locally Lipschitz continuous and 00 is a regular value of each $\Delta_{i}$, then the trajectory of the augmented system (24) under policy $(\boldsymbol{\mathbf{u}},\boldsymbol{\mathbf{w}})=\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$ exists, is unique, and satisfies $\boldsymbol{\mathbf{x}}(t)\in\mathcal{X}$ and $\boldsymbol{\mathbf{v}}(t)\in\mathcal{V}$ for all $t\geq 0$, given $\big{(}\boldsymbol{\mathbf{x}}(0),\boldsymbol{\mathbf{v}}(0)\big{)}\in\tilde{\mathcal{C}}$.
Proof 3.13.
Feasibility of the DSM-CBF policy follows by Corollary 3.8. Any locally Lipschitz controller in $\tilde{\mathcal{K}}_{\alpha}$ certifies the local (in time) existence and uniqueness of the solution to the closed-loop system [35]. Due to the the invariance of $\tilde{\mathcal{C}}$, there can be no finite-time escape instances and the existence and uniqueness holds for all $t\geq 0$. Safety follows from $\tilde{\mathcal{C}}\subset\mathcal{X}\times\mathcal{V}$.
@@
Convexity of the set $\tilde{\mathcal{K}}_{\alpha}(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})$, and therefore problem (31), is linked to convexity of $\mathcal{U}$. In particular, if the input constraint set $\mathcal{U}$ is polyhedral, the optimization problem (31) becomes a quadratic program (QP) for all $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}}$, which can be solved efficiently.
@@
3.3 Existence and Uniqueness of Solutions
Given the prevalence of optimization-based policies in constrained control approaches such as MPC and CBFs, numerous works have studied their regularity properties [37, 12, 38].
Given any initial condition $(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}})\in\tilde{\mathcal{C}},$ the optimization-based policy $\beta$ for DSM-based CBFs, defined in (31), is always feasible. If, additionally, it is locally Lipschitz continuous, then the trajectory for the closed-loop system
$\begin{bmatrix}\dot{\boldsymbol{\mathbf{x}}}\\ \dot{\boldsymbol{\mathbf{v}}}\end{bmatrix}=\begin{bmatrix}f(\boldsymbol{\mathbf{x}})\\ 0\end{bmatrix}+\begin{bmatrix}g(\boldsymbol{\mathbf{x}})&0\\ 0&I\end{bmatrix}\beta(\boldsymbol{\mathbf{x}},\boldsymbol{\mathbf{v}}),$ (32)